• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 22:37:45 on June24,2004

    Entry number 126734

    This entry is a followup to: 126708

    keyword=prof. scanner, part 3

    HV on scanner PMT set to 1650.
    Run 1169 -- brief, wrong prescales.
    (T5 at 690 kHz, T6 at 82 kHz)
    Run 1170 better prescales.

    Fig 1 shows the ADC spectra for 1650 V.

    Conditions otherwise same as before, P = -10%.
    scanner at X=17, Y=33 on both arms.
    Piotr says their Y is my X. But I write what LabView says.
    Current was 0.5 uA

    FIGURE 1